Solve the following equation. Show all your work.
logx(10-3x)=2

Answer by nerdybill(5404)

logx(10-3x)=2
.
If you meant "log base x" then:
logx(10-3x)=2
10-3x=x^2
0 = x^2+3x-10
0 = (x+5)(x-2)
.
x = {-5, 2}